Bubbly:

Identify the numbers that are irrational. Select the two answers that apply. Group of answer choices 2 squared 0.8333... 3pi/pi 0.4 pi 4 squared

Bubbly:

\[A.\sqrt{2}\]

Bubbly:

\[C. \frac{ 3\Pi}{ \Pi } \]

Bubbly:

\[E.\Pi \]

Bubbly:

\[F.\sqrt{4}\]